Let X = {a, b, c} defined by f : X
1
Expert's answer
2021-09-16T00:43:50-0400

Since, we have incomplete data, we assume as follows:

Let X = {a, b, c} defined by f : X"\rightarrow" R such that f(x) = x+1 "\forall x\in X" . Find all such values of f(x).

Solution:

f(x) = x+1

f(a) = a+1

f(b) = b+1

f(c) = c+1

Thus, all values of f(x) are {a+1, b+1, c+1}.